A bat hits a ball of 0.145 kg. Initially, the ball
Answer #1

According to the given problem,

A) Impluse = Δp = mv2 - mv1

x- component :

Jx = 0.145(-65m/s*cos30º - 60m/s) = -16.8623 kg-m/s

y- component:

Jy =  0.145kg * 65m/s * sin30º = 4.7125 kg-m/s

J = 17.5 kg-m/s

B) Force :

Fy = mΔv/Δt = 0.145kg * 65m/s * sin30º / 0.00175s = 2692.857 N
horizontal:
Fx = 0.145kg * (-65m/s*cos30º - 60m/s) / 0.00175s = -9635.588 N

Favg = 10004.8 N